# A Lot of Talk and a Badge: An Exploratory Analysis of Personal Achievements in GitHub

### 30 Dec 2023 v2

@@ 5.4.2 Community discussions analysis
Overall, we mined all the Q&A threads related to personal badges from June (i.e., the launch of the new feature) to December 2022. We retrieved 312 questions, containing 937 answers and 781 replies. Subsequently, we performed a thematic analysis of the questions we had retrieved.
@@
To carry out the thematic analysis, we divided the entire set of 312 posts into subsets of about 100 elements; then, the first two authors independently analyzed the first subset. We identified common and different categories. The differences were resolved through discussions with the entire team. After completing the analysis of the first subset, we obtained an initial coding schema, which was applied independently to the second subset of posts. A similar approach was followed to resolve differences and update the coding schema with new or refined categories. When necessary, the changes were propagated back to the previous subset. After resolving all the differences, we moved on to the third (final) subset, replicating the same approach. We used Taguette777https://www.taguette.org to support the annotation process and a spreadsheet to define the coding schema.
@@
Thematic analysis revealed four main themes (see Table 4): Info requests, Feedback, Improvements, and Other. In the following paragraphs, we describe each theme and the corresponding codes, exemplifying the most relevant concepts with excerpts from the discussion threads.
@@
The prevalent theme that emerged from our analysis is the request for information about personal badges – namely, Info requests. This theme encompasses 122 distinct questions, a number that doubles those recorded for the next two main themes, i.e., Improvements and Feedback. Arguably, the prevalence of questions from the theme Info requests can be explained by the choice of GitHub not to disclose explicit information on personal badges; this choice has led several users of the platform to seek unofficial information on the community forum. Within this theme, the most frequent code is How to get a badge, assigned to 79 questions regarding the requirements needed to earn a specific badge (e.g., “How do you get the Galaxy Brain Badge?”). Similarly, 17 more questions were coded as How to get badges; these represent requests for help on how to earn more badges in general (“How to get a new badge on GitHub?”). The remaining 26 questions from this theme were coded as General info; these are generic requests for information about the badges feature, the most common being the full list of the currently available personal badges in GitHub (e.g., “How can I see all available badges in GitHub? Please Help!!”).
@@
The next major theme identified is the Improvements category, which includes 60 questions and 2 codes: Badge(s) proposals and UX suggestions. Under Badge(s) proposals, we collected 22 questions that represent requests for the addition of new personal badges. For instance, three users recommended adding an additional badge to reward code-review activity (e.g., “I’ve learned a lot from code reviewing others’ code and being code reviewed by others”); furthermore, three other users proposed the addition of organization-specific badges (e.g., “to celebrate the first PR merge of a new employee”). Some users even suggested creative names for their new badge ideas, such as the “Issue Muncher” badge—keeping track of “how many issues one has helped to close through linked PRs”—and the “Bug Hunter” badge—rewarding reports of bugs in GitHub. Also, a user suggested that the personal badge feature be extended to showcase actual professional achievements, such as the Linux Foundation Certified Kubernetes Administrator certificate or the Google Certification Academy certificate. Finally, a couple of more questions with this code were concerned about how to suggest new types of badge, showing that some users are willing to participate in the definition of future personal achievements. Regarding the UX suggestions code, we identified 40 threads. In particular, 21 questions concerned a feature request, the most popular (with 8 questions) being the possibility to selectively hide/show the earned badges on user profile pages; other recurring feature requests are: enabling custom badge ordering in profiles (2), taking historical data into account when defining badge unlocking rules (2), and enabling a summary view displaying all possible achievements (2). Another common type of UX-related question (11 questions) represents negative feedback on the graphical appearance of badges (e.g., “These new badges are too cartoonish”, “The designs could be more professional”) or suggestions on how to improve the display of badges (e.g., “Make the old badges look like the new ones.”)
@@
The most significant theme in relation to RQ4 is Feedback, encompassing the codes Positive, Negative (in general), and Negative (specific), for a total of 53 questions. Although some questions (12) report positive feedback, in the form of short and generic statements of appreciation for the feature of personal badges, most of the questions collected within this category were coded as negative feedback, either generic (25 questions) or feature-specific (16 questions). Regarding generic negative feedback, several users reported disregarding the new feature, and 17 out of 25 asked how to opt out (or declared their intention to do so). Frequent causes for the negative feedback were: i) aversion to the gamification of a professional environment like GitHub and the “childish” style of badges – e.g., “[they] look like Candy Crash Saga. These weird cartoon emoji are massive, and distract the eye from actual content” (11 questions); ii) skepticism about the ability of personal badges to convey developers’ skills – e.g., “unnecessary gamification that doesn’t help assess the performance/dedication/talent/achievement of a developer” (2 questions); iii) concerns about the potential negative effects of badges on user behaviors – e.g., “PRs created not for the purpose of making a meaningful contribution, but simply to get another badge checked off their list.”, “the achievements system sets up terrible incentives for folks to create PRs, not for the purpose of making a meaningful contribution, but simply to get another badge checked off their list.” – and the whole community – e.g., “[Badges] risk becoming a driver for ivory tower superiority and community tribalism, which is something the developer community already struggles with”, “I’m not listening to your feature request because I have badges and you don’t” (5 questions).
Regarding badge-specific negative feedback, most complaints related to the badge YOLO (7 questions), which is perceived by several developers as a source of shame and an unfair mark to obtain, especially if earned within single-person projects (e.g., “I feel having the YOLO badge does not send the right message, especially to recruiters”). Another common criticism (4 questions) concerned the unclear requirements for unlocking the Pull Shark badge, sometimes perceived as biased (e.g., “Pull Shark achievement should work [only] for repositories I don’t own. […] we will always merge our own pull requests, but not everyone will merge our pull request and that’s the point”).
Finally, the theme Other gathered 77 questions, mostly coded as requests of support for a Technical issue (e.g., Pull Shark shows the wrong merged PR number for me”), and 4 questions coded as Miscellaneous. In addition, we excluded 31 questions when their content was deemed off-topic or when the text body was empty, not in English, or incomprehensible.
@@ 6 Discussion
Research questions. We answered four research questions. First (RQ), we explored the diffusion of badges among a sample of 6,022 GitHub developers as of December 2022 (Table 1) and their evolution over six months, since their introduction in June 2022 (Figure 2). We found that all developers except one own at least one badge and that the number of those who opted out of displaying badges and chose to make their profile pages private has been steadily increasing. Then (RQ), we investigated whether owning badges is associated with signaling certain hypothesized developers’ qualities and dispositions. The initial results of our exploratory analysis (Figure 3) showed only a large difference for developers owning the badge Starstruck, gained by those who get their repositories starred and a signal of increased popularity. Furthermore, the regression analysis (RQ) showed that the only effect of implementing personal badges in GitHub has been a large increase in the number of followers of developers who own the Starstruck badge (Table 3). Finally (RQ), the analysis of the survey responses showed several shortcomings in the current implementation of badges, such as the limited types of badges currently existing and the lack of those more accurately reflecting contributions to projects and years of experience. Besides, the analysis of community discussions about personal badges revealed that GitHub users are willing to know more about badges and the requirements set out for earning them; also, part of the community is willing to participate in the definition of future achievements. However, the discussions analyzed showed the clear prevalence of negative opinions on badges, which confirms the skepticism of several users about their ability to adequately convey the skills of a developer and the concerns of the community about the potential drawbacks of gamifying GitHub.
@@
Badges as signals. The results of the user survey confirmed that GitHub developers do look at profile pages and are aware of the elements therein—so much so that several even decided to opt out of displaying badges. As such, these pages have the potential to act as hubs that make socially translucent signals visible and readily available, and support awareness of collaborators’ behavior (Erickson and Kellogg, 2000). This result is consistent with those reported by Shami et al. (2009) who used the signaling theory as a conceptual framework to investigate how users of online social platforms rely on digital artifacts for ‘people sensemaking,’ i.e., use portions of profile pages as proxies used to infer unknown coworkers’ expertise.
@@
However, the visibility of personal badges is not sufficient to ensure their reliability. According to signaling theory (Donath, 2007), to be reliable, assessment signals must be both observable and costly to produce. Our findings suggest that most of the currently available personal badges, although they reflect the properties possessed, do not send reliable assessment signals because they may not be costly enough to produce. The only exception to general unreliability is the Starstruck badge signaling popularity. One potential explanation for this finding is that the Starstruck badge is awarded to a developer as a result of the interest of others (stars) in their repositories: creating a repository implies an effort broader in scope, if not harder, as compared to other badges, which require the developers to take one action (e.g., answer a question, merge a pull request). This result is somewhat consistent with those reported by Trockman et al. (2018) who found that popularity-related repository badges are more reliable as assessment signals. One major difference with their work is that repository badges are chosen—maintainers select what they intend to signal by adding them to README pages (e.g., the count of downloads badge to signal popularity, up-to-date dependency badge to signal the lack of known security risks). Conversely, personal badges are a truer gamification element—they are unlocked and displayed on personal pages—and the only control developers have over them is choosing whether to opt out of showing them altogether.
@@
Badges as a gamification mechanism. According to Hunter and Werbach (2012), users of a gamified environment go through a journey, from onboarding to scaffolding and, finally, mastery. As argued by Sasso et al. (2017), implementing game mechanisms such as badges to gamify a software development environment is an iterative and far from trivial undertaking. Our analyses suggest that the leveling process of personal badges in GitHub needs to be adjusted, especially during the onboarding and mastery stages. On the one hand, some junior developers among the survey respondents reported that it “felt nice” to discover the first badge on their profile. However, the easier-to-achieve onboarding badges (i.e., YOLO and Quickdraw) are obtained through practices that are not to be encouraged in software development (i.e., closing issues and PRs quickly or without code review), and in the community discussions some users referred to these badges as “shameful” and “trivial.” Conversely, more senior developers argued that they feel that current badges are “too easily achieved, not like in Stack Overflow” and complained about the lack of badges reflecting their years of experience and project contributions (“They don’t give the impression that I’m a professional developer, but that I’m a novice coder who’s just finished my first project.”).
@@
In our study, we found an increasing number of developers (up to 23%) who made their profile page private or opted out of displaying personal badges. Albeit neither the survey or the discussion analyses helped us to understand the motivations, one possible explanation is that GitHub developers are conscious that the elements on their profile page may influence the processes of impression formation (Marlow et al., 2013; Singer et al., 2013) and people sensemaking (Shami et al., 2009). For instance, more and more employers look at information gleaned from social networking sites during the hiring process to assess job applicants (car, 2017), and their social media image has an influence on the chance of being hired (Bohnert and Ross, 2010). However, Marlow and Dabbish (2013) found that, albeit conscious, only few GitHub developers engaged in efforts to clean up their profile page because deemed too costly. Our findings show that GitHub developers may have grown even more conscious over time and that the decision to allow them to easily opt out of displaying personal badges is a step in the right direction. Nonetheless, GitHub platform designers should consider implementing the ability to selectively hide/show badges on user profile pages – a popular feature requests mentioned by several survey participants; this might alleviate the discontent caused by badges like YOLO and Quickdraw, which mark and publicly expose undesirable and unprofessional user behaviors.
@@
In terms of usability, to meet user expectations, platform designers should also consider making the progress toward the achievement of badges more explicit, especially in the case of debated badge requirements (e.g., those of the Pull Shark badge), and graphic designers should also revise the appearance of badges to give them a more professional and consistent ‘look and feel.’
@@
Furthermore, the current implementation of badges in GitHub appears underdeveloped. First, as discussed earlier, the current number of badges and the implementation of the leveling system do not seem to fit well with the onboarding and mastery stages typical of a gamified environment. As such, we suggest platform designers consider adding more: i) onboarding badges, that are appealing to ‘newcomers’ (e.g., junior developers) without fostering bad software development practices such as no or quick code reviews; ii) mastery badges, that reflect long-running, broad-in-scope achievements that are appealing to and feel ‘earned’ by more experienced, senior developers. Second, platform designers should consider rolling out more of them all at once rather than slowly, as some users complained that they “all have almost the same badges.” Sarma et al. (2016) developed Visual Resume, a tool that aggregates activity traces collected from GitHub and Stack Overflow to aid employers in the hiring process; they evaluated the tool with managers and other technical personnel and found that they prefer cues in aggregated form (e.g., activity summaries) and community-generated endorsements (e.g., reputation) to inform hiring decisions and form impressions. Their findings may provide platform designers with useful ideas to implement more reliable personal badges, as participants in that study expressed a desire to have quick access to features indicating whether commits are central to the code base and whether they make significant changes rather than tweak/fix code. Third, platform designers should further reflect on the unintended side effects of the decision not to provide any official documentation regarding badges. Our findings show that this decision implies that no clear indication is given of what cues the badges are intended to signal beyond their name and image, with a potential impact on impression formation and people sensemaking, for example, in collaborative development (Marlow et al., 2013; Singer et al., 2013). As of this writing, two more badges have been discovered, namely Heart On Your Sleeve and Open Sourcerer but the community still has not confirmed what they mean or how to achieve them.888https://github.com/drknzz/GitHub-Achievements999https://github.com/github-profile-achievements/english
@@
Finally, we urge platform designers to reflect about the potential side effects that displaying personal badges may have on the hiring process, as revealed by our analysis of community discussions (“A potential employer glancing at my profile, perhaps unfamiliar with this very niche GitHub-specific feature, might assume a negative impression of me”). Albeit GitHub is designed as a platform to support collaborative software development, employers and job seekers also use it as a recruitment tool (Capiluppi et al., 2013; Greene and Fischer, 2016). Marlow and Dabbish (2013) analyzed the signals given off by activity traces displayed on GitHub profile pages through the lens of signaling theory. After interviewing managers and developers who use GitHub as part of the hiring and job application process, they found that interviewees valued the insights provided by GitHub accounts as more reliable and verifiable than a static list of individuals’ skills. In fact, according to Walther et al. (2009), when forming an impression, information provided by third parties (i.e., GitHub in this case), is more reliable than self reports. Nonetheless, Marlow and Dabbish found that the reliability of the activity traces as signals varied and their use was directly related to the evaluation cost. In particular, they found that activity traces send reliable assessment signals when they reflect (i) history of activity (e.g., commitment to an open source project over years) and (ii) networking (e.g., contributing commits to high-status OSS projects). Still, employers only chose to look at those activity traces that are easy to verify quickly. Hence, consistently with this evidence and the signaling theory, we suggest that platform designers overhaul personal badges so that they provide quick access to aggregated traces of activity history and collaboration in high-status projects, which are costly to produce for job seekers and, at the same time, affordable for employers and fellow developers to verify.

### 02 Feb 2024 v3

@@ 5.4.2 Community discussions analysis
Overall, we extracted all the Q&A threads related to personal badges from June (i.e., the launch of the new feature) to December 2022 and retrieved 312 questions.
@@
Subsequently, we performed a thematic analysis. We divided the entire set of 312 posts into three subsets of approximately 100 elements; then, the first two authors independently analyzed the first subset. We identified common and different categories. The differences were resolved through discussions with the entire team. For the two other subsets, we followed the same approach, with changes to the coding schema propagated back to the previous batche(s) if necessary.
@@
Thematic analysis revealed four main themes (see Table 4): Info requests, Feedback, Improvements, and Other. Next, we describe each theme and the corresponding codes, exemplifying the most relevant concepts with excerpts.
@@
The prevalent theme that emerged from our analysis is the request for information about personal badges – namely, Info requests. This theme encompasses 122 distinct questions, almost twice as many as the other two categories. Arguably, the prevalence of questions on this theme is explained by the choice of GitHub not to disclose explicit information on personal badges; this choice has led several users to seek unofficial information on the community forum. Within Info requests, the most frequent code is How to get a badge, assigned to 79 questions regarding the requirements needed to earn a specific badge (e.g., “How do you get the Galaxy Brain Badge?”). Similarly, 17 more questions were coded as How to get badges; these represent requests for help on how to earn more badges in general (“How to get a new badge on GitHub?”). The remaining 26 questions from this theme were coded as General info; these are generic requests for information about the badges feature, the most common being the full list of the currently available personal badges in GitHub (e.g., “How can I see all available badges in GitHub? Please Help!!”).
@@
The next identified major theme is the Improvements category, which includes 60 questions and two codes. Under Badge(s) proposals, we collected 22 questions that represent requests for the addition of new personal badges. For example, three users recommended adding an additional badge to reward code-review activity (e.g., “I’ve learned a lot from code reviewing others’ code and being code reviewed by others”); furthermore, three other users proposed the addition of organization-specific badges (e.g., “to celebrate the first PR merge of a new employee”). Some users even suggested creative names for their new badge ideas, such as the “Issue Muncher” badge, to keep track of “how many issues one has helped close through linked PRs”, and the “Bug Hunter” badge, to reward bug reporting. Also, a user suggested that the personal badge feature be extended to showcase actual professional achievements, such as the Linux Foundation Certified Kubernetes Administrator and Google Certification Academy certificates. Finally, a couple of additional questions with this code asked how to suggest new types of badge, showing that some users are willing to participate in the definition of future personal achievements. Regarding the UX suggestion code, we identified 40 threads. In particular, 21 questions concerned a feature request, the most popular (8 questions) being the possibility to selectively hide/show the earned badges on user profile pages; other recurring feature requests are: enabling custom badge ordering in profiles (2), taking historical data into account when defining badge unlocking rules (2), and enabling a summary view displaying all possible achievements (2). Several UX-related questions (11) were posted to criticize the graphical appearance of badges (e.g., “These new badges are too cartoonish”, “The designs could be more professional”) or suggest how to improve it (e.g., “Make the old badges look like the new ones”).
@@
The most significant theme in relation to RQ4 is Feedback, which includes the codes Positive, Negative (in general), and Negative (specific), for a total of 53 questions. Although some questions (12) report positive feedback, in the form of short and generic statements of appreciation for the feature of personal badges, most of the questions collected within this category were coded as negative feedback, either generic (25 questions) or feature-specific (16 questions). Regarding generic negative feedback, several users reported disregarding the new feature, and 17 out of 25 asked how to opt out. Frequent causes for the negative feedback were: i) aversion to the gamification of a professional environment like GitHub and the “childish” style of badges (11 questions) – e.g., “[they] look like Candy Crash Saga. These weird cartoon emoji are massive, and distract the eye from actual content”; ii) skepticism about the ability of personal badges to convey developers’ skills (2 questions) – e.g., “unnecessary gamification that doesn’t help assess the performance/dedication/talent/achievement of a developer”; iii) concerns about the potential negative effects of badges on user behaviors (5 questions) – e.g., “PRs created not for the purpose of making a meaningful contribution, but simply to get another badge checked off their list.”, “[achievements] set up terrible incentives for folks to create PRs, not for the purpose of making a meaningful contribution, but simply to get another badge checked off their list” – and the whole community – e.g., “[Badges] risk becoming a driver for ivory tower superiority and community tribalism, which is something the developer community already struggles with”, “I’m not listening to your feature request because I have badges and you don’t”.
Regarding badge-specific negative feedback, most complaints related to the badge YOLO (7 questions), which is perceived by several developers as a source of shame and an unfair mark to obtain, especially if earned within single-person projects (e.g., “I feel having the YOLO badge does not send the right message, especially to recruiters”). Another common criticism (4 questions) concerned the unclear requirements for unlocking the Pull Shark badge, sometimes perceived as biased (e.g., “Pull Shark achievement should work [only] for repositories I don’t own. […] we will always merge our own pull requests, but not everyone will merge our pull requests”).
Finally, the theme Other gathered 77 questions, mostly coded as requests of support for a Technical issue (e.g., Pull Shark shows the wrong merged PR number for me”), and 4 questions coded as Miscellaneous. In addition, we excluded 31 questions when their content was empty, deemed off topic, or not in English.
@@ 6 Discussion
Research questions. We answered four research questions. First (RQ), we explored the diffusion of badges among a sample of 6,022 GitHub developers as of December 2022 (Table 1) and their evolution over six months, since their introduction in June 2022 (Figure 2). We found that all developers except one own at least one badge and that the number of those who opted out of displaying badges and chose to make their profile pages private has been steadily increasing. Then (RQ), we investigated whether the ownership of badges is associated with signaling certain hypothesized developers’ qualities and dispositions. The initial results of our exploratory analysis (Figure 3) showed only a large difference for developers owning the badge Starstruck, gained by those who get their repositories starred and a signal of increased popularity. Furthermore, regression analysis (RQ) showed that the only effect of implementing personal badges in GitHub has been a large increase in the number of followers of developers who own the Starstruck badge (Table 3). Finally (RQ), the analysis of the survey responses showed several shortcomings in the current implementation of badges, such as the limited types of badges currently existing and the lack of those more accurately reflecting contributions to projects and years of experience. Besides, the analysis of community discussions about personal badges revealed that GitHub users are willing to know more about badges and the requirements set out for earning them; also, part of the community is willing to participate in the definition of future achievements. However, the discussions analyzed showed a clear prevalence of negative opinions on badges, highlighting the skepticism of the community about their ability to adequately convey developers’ skills and a general concern about the potential drawbacks of GitHub gamification.
@@
Badges as signals. The results of the user survey confirmed that GitHub developers do look at profile pages and are aware of the elements therein—so much so that several even decided to opt out of displaying badges. As such, these pages have the potential to act as hubs that make socially translucent signals visible and readily available, and support awareness of collaborators’ behavior (Erickson and Kellogg, 2000). This result is consistent with those reported by Shami et al. (2009) who used the signaling theory (Donath, 2007) as a conceptual framework to investigate how users of online social platforms rely on digital artifacts for people sensemaking, i.e., use portions of profile pages as proxies used to infer unknown coworkers’ expertise.
@@
However, the visibility of personal badges is not sufficient to ensure their reliability. According to signaling theory, to be reliable, assessment signals must be both observable and costly to produce. Our findings suggest that most of the currently available personal badges, although they reflect the properties possessed, do not send reliable assessment signals because they may not be costly enough to produce. The only exception to general unreliability is the Starstruck badge signaling popularity. One potential explanation for this finding is that the badge is awarded to a developer as a result of the interest of others (stars) in their repositories: creating a repository implies an effort broader in scope, if not harder, as compared to other badges, which require the developers to take one action (e.g., answer a question, merge a pull request). This result is somewhat consistent with those reported by Trockman et al. (2018), who found that popularity-related repository badges are more reliable as assessment signals. One major difference with their work is that repository badges are chosen—maintainers select what they intend to signal by adding them to README pages (e.g., the count of downloads to signal popularity, up-to-date dependencies to signal security). Conversely, personal badges are a truer gamification element – they are unlocked and displayed on personal pages – and the only control developers have over them is to not show them altogether.
@@
Badges as a gamification mechanism. According to Hunter and Werbach (2012), users of a gamified environment go through a journey, from onboarding to scaffolding and, finally, mastery. Implementing game mechanisms such as badges to gamify a software development environment is an iterative and far from trivial undertaking (Sasso et al., 2017). Our analyses suggest that the leveling process of personal badges in GitHub needs to be adjusted, especially during the onboarding and mastery stages. On the one hand, some junior developers among the survey respondents reported that it “felt nice” to discover the first badge on their profile. However, the easier-to-achieve onboarding badges (i.e., YOLO and Quickdraw) are obtained through practices that are not to be encouraged in software development (i.e., closing issues and PRs quickly or without code review) – in the community discussions, some users referred to these badges as “shameful” and “trivial.” In contrast, more senior developers argued that they feel that current badges are “too easily achieved, not like in Stack Overflow” and complained about the lack of badges reflecting their years of experience and project contributions (“They don’t give the impression that I’m a professional developer, but that I’m a novice coder who’s just finished my first project.”).
@@
In our study, we found an increasing number of developers (up to 23%) who made their profile page private or opted out of displaying personal badges. Although neither the survey nor the discussion analyses helped us to understand the motivations, one possible explanation is that GitHub developers are aware that the elements on their profile page may influence the processes of impression formation (Marlow et al., 2013; Singer et al., 2013) and people sensemaking (Shami et al., 2009). For example, more and more employers look at the information gathered from social networking sites during the hiring process to assess job applicants (car, 2017), and their image on social media has an influence on the chance of being hired (Bohnert and Ross, 2010). However, previous research found that, although conscious, only few GitHub developers engaged in efforts to clean up their profile page because deemed too costly (Marlow and Dabbish, 2013). Our findings show that GitHub developers may have grown more aware over time and that the decision to allow them to easily opt out of displaying personal badges is a step in the right direction. Nevertheless, GitHub platform designers should also consider implementing the ability to selectively hide/show badges on user profile pages – a feature requested by several survey participants; this might alleviate the discontent caused by badges like YOLO and Quickdraw, which mark and publicly expose undesirable and unprofessional user behaviors.
@@
In terms of usability, platform designers should also consider making the progress toward badge achievement more explicit, especially in the case of debated badge requirements (e.g., those of the Pull Shark badge), and graphic designers should also revise the appearance of badges to give them a more professional and consistent ‘look and feel.’
@@
Furthermore, the current implementation of badges in GitHub appears underdeveloped. First, as discussed earlier, the current number of badges and the implementation of the leveling system do not seem to fit well with the onboarding and mastery stages typical of a gamified environment. As such, we suggest platform designers consider adding more: i) onboarding badges, that are appealing to newcomers (e.g., junior developers) without fostering bad software development practices such as no or quick code reviews; ii) mastery badges, that reflect long-running, broad-in-scope achievements that are appealing to and feel ‘earned’ by more experienced, senior developers. Second, platform designers should consider rolling out more of them all at once rather than slowly, as some users complained that they “all have almost the same badges.” Sarma et al. (2016) developed Visual Resume, a tool that aggregates activity traces collected from GitHub and SO to aid employers in the hiring process; they found that managers and other technical personnel prefer cues in aggregated form (e.g., activity summaries) and community-generated endorsements (e.g., reputation) to inform hiring decisions and form impressions. Their findings may provide platform designers with useful ideas to implement more reliable personal badges, such as having quick access to features indicating whether commits are central to the code base and whether they make significant changes rather than tweak/fix code. Third, platform designers should further reflect on the unintended side effects of the decision not to provide any official documentation regarding badges. Our findings show that this decision implies that no clear indication is given of what cues badges are intended to signal beyond their name and image, with a possible impact on impression formation and people sensemaking during collaborative development (Marlow et al., 2013; Singer et al., 2013). As of this writing, two more badges have been discovered, namely Heart On Your Sleeve and Open Sourcerer, but the community still has not confirmed what they mean or how to achieve them.777https://github.com/drknzz/GitHub-Achievements888https://github.com/github-profile-achievements/english
@@
Finally, we urge platform designers to reflect on the potential side effects that displaying personal badges may have on the hiring process, as revealed by our analysis of community discussions (“A potential employer glancing at my profile, perhaps unfamiliar with this […] feature, might assume a negative impression of me”). Although GitHub is designed as a platform to support collaborative software development, employers and job seekers also use it as a recruitment tool (Capiluppi et al., 2013; Greene and Fischer, 2016). Marlow and Dabbish (2013) analyzed the signals given off by the activity traces displayed on GitHub profile pages through the lens of signaling theory. After interviewing managers and developers who use GitHub as part of the hiring and job application process, they found that interviewees valued the insights provided by GitHub accounts as more reliable and verifiable than a static list of individuals’ skills. In fact, according to Walther et al. (2009), when forming an impression, information provided by third parties (i.e., GitHub in this case), is more reliable than self reports. However, Marlow and Dabbish found that the reliability of the activity traces as signals varied and that their use was directly related to the evaluation cost. In particular, they found that activity traces send reliable assessment signals when they reflect the (i) history of activity (e.g., commitment to an open source project over years) and (ii) networking (e.g., contributing to high-status OSS projects). Still, employers only chose to look at those activity traces that are easy to verify quickly. Hence, consistently with this evidence and the signaling theory, we suggest that platform designers overhaul personal badges so that they provide quick access to aggregated traces of activity history and collaboration in high-status projects, which are costly to produce for job seekers and, at the same time, affordable for employers to verify.